# Communication-Efficient Stochastic Distributed Learning

### v1

@@ C-A Key bounds
$\delta\coloneqq 1-\frac{|\tilde{\lambda}_{\max}|\rho\tau\gamma}{2},\tilde{c}_{0}\coloneqq\gamma^{2}\tau\tilde{\beta}_{0},\tilde{c}_{1}\coloneqq\gamma^{3}\tau^{2}\tilde{\beta}_{1},$
$\tilde{c}_{2}\coloneqq{12L^{2}\gamma^{2}}\|\widehat{\mathbf{V}}^{-1}\|^{2}N,\tilde{c}_{3}\coloneqq\gamma^{2}\tau\tilde{\beta}_{3},$
$\tilde{q}_{1}:=\tilde{c}_{0}\gamma\tau\kappa_{1}+\tilde{c}_{1},\tilde{q}_{2}\coloneqq\tilde{c}_{2},\quad\tilde{q}_{3}:=16\tau^{3}\gamma^{2}N\tilde{c}_{0},$
$\tilde{q}_{4}:=\left(48\tau^{2}\gamma^{2}L^{2}\tilde{c}_{0}+\tilde{c}_{3}\right)/(1-\delta).$
Proof.
When $\gamma<\min\{\bar{\gamma}_{1},\bar{\gamma}_{11}\}$ we have
$\gamma\tau\leq\frac{2}{|\tilde{\lambda}_{\max}|\rho},\quad|\tilde{\lambda}_{\min}|\rho\tau\gamma<2,$ (59)
$6\gamma^{2}\tau(3L^{2}+\rho^{2}d_{u}^{2})<\frac{1/4}{\tau-1}.$
Substituting (48) into (28) then yields
$\|\widehat{\mathbf{h}}_{k}\|^{2}\leq\tilde{c}_{0}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\tilde{c}_{1}\|\widehat{\mathbf{d}}_{k}\|^{2}+\tilde{c}_{2}\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}+\tilde{c}_{3}t_{k},$
and together with (49) and (17), it proves that (58) holds. ∎
C-B Theorem 2
Based on (30), substituting $y=\bar{x}_{k+1}$ and $z=\bar{x}_{k}$ into (37) and using (27), (48), we get
@@ C-B Theorem 2
$\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\gamma\mathbb{E}[\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}(\sum_{t}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)+\epsilon_{k})\rangle]$
$+2\gamma^{2}\tau L\mathbb{E}[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\frac{2\gamma^{2}\tau L}{N}(2L^{2}\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]+2L^{2}\mathbb{E}[{t}_{k}])$
$+\frac{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau(\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]+\tau\mathbb{E}[\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}])}{N}.$
Similarly to Appendix B-B, using $2\langle a,b\rangle=\|a\|^{2}+\|b\|^{2}-\|a-b\|^{2}$ we have
-- $\mathbb{E}[\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}\sum_{t}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)+\frac{\epsilon_{k}}{N}\rangle]$
$\leq-\frac{\tau}{4}\mathbb{E}[\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}]-\frac{1}{2}\mathbb{E}[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\left(\frac{L^{2}}{2N}+\frac{2\rho^{2}d_{u}^{2}}{N}\right)\mathbb{E}[\widehat{\mathbf{\Phi}}_{k}\|^{2}]+\frac{2\rho^{2}d_{u}^{2}\tau}{N}\mathbb{E}[\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}].$
Combining the two inequalities above, and using (16), yields
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\frac{\gamma\tau}{4}\mathbb{E}[\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}]$ (60)
$-\frac{\gamma}{2}(1-4\gamma\tau L)\mathbb{E}\left[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}\right]$
$+\left(\frac{\gamma L^{2}}{2N}+\frac{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2}+4\gamma^{2}\tau L^{3}}{N}\right)\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]$
$+\frac{18\gamma\rho^{2}d_{u}^{2}\tau+18\gamma^{2}\rho^{2}d_{u}^{2}\tau^{2}L}{N}\frac{2\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+\frac{4\gamma^{2}\tau L^{3}\mathbb{E}[t_{k}]}{N}.$
Combining (60) together with (49) and (53), and using $\gamma\leq 1$, yields
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\frac{\gamma\tau}{4}\mathbb{E}[\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}]$
$-\frac{\gamma}{2}(1-4\gamma\tau L)\mathbb{E}[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\left(k_{0}k_{2}+2(k_{3}k_{0}+k_{4})s_{2}\right)\mathbb{E}[\|\nabla F(x)\|^{2}]$
$+\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}],$
@@
$k_{0}\coloneqq\alpha_{0}\gamma,k_{1}\coloneqq\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2}$
$k_{2}\coloneqq 16\tau^{3}\gamma^{2}N,k_{3}\coloneqq 48\tau^{2}\gamma^{2}L^{2}$
$k_{4}\coloneqq\frac{4\gamma^{2}\tau L^{3}}{N},k_{5}\coloneqq\frac{36\gamma^{2}\rho d_{u}^{2}\tau^{2}+36\gamma^{3}\rho d_{u}^{2}\tau^{3}L}{N|\tilde{\lambda}_{\max}|}.$
Letting $\gamma\leq\min\{1,\bar{\gamma}_{12},\bar{\gamma}_{13}\}$, then
$k_{0}k_{2}+2(k_{3}k_{0}+k_{4})s_{2}\leq\frac{\gamma\tau}{8},\quad 4\gamma\tau L\leq\frac{3}{4},$ (61)
@@
$-\frac{\gamma}{8}\mathbb{E}[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}].$
@@
$\mathcal{D}_{k}$ $\leq\frac{8}{\gamma\tau}(\mathbb{E}[\tilde{F}\left(\bar{x}_{k}\right)]-\mathbb{E}[\tilde{F}\left(\bar{x}_{k+1}\right)])$ (62)
$+\frac{8}{\gamma\tau}\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}],$
where $\mathcal{D}_{k}$ is defined in (7), and $\tilde{F}\left(\bar{x}_{k}\right)=F\left(\bar{x}_{k}\right)-F(x^{*})$. Summing (62) over $k=0,1,\ldots,K-1$ and using $-\tilde{F}\left(\bar{x}_{k}\right)\leq 0$, it holds that
$\sum_{r=0}^{K-1}\mathcal{D}_{k}\leq\frac{8\tilde{F}\left(\bar{x}_{0}\right)}{\gamma\tau}$ (63)
$+\frac{8\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)}{\gamma\tau}\sum_{k=0}^{K-1}\|\mathbb{E}[\hat{\mathbf{d}}_{k}\|^{2}].$
According to (58) and (53), we derive that $\forall k\geq 0$,
$\mathbb{E}[\|\widehat{\mathbf{d}}_{k+1}\|^{2}]\leq\left(\delta+\frac{\tilde{q}_{1}}{1-\delta}+2\tilde{q}_{4}(s_{0}+s_{1})\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$ (64)
$+\frac{\tilde{q}_{2}}{1-\delta}\mathbb{E}[\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}]+\left(\frac{\tilde{q}_{3}}{1-\delta}+2\tilde{q}_{4}s_{2}\right)\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]$
@@
$\tilde{R}=\max\left\{\frac{\tilde{q}_{2}\tau^{2}}{1-\delta},\frac{\tilde{q}_{3}}{1-\delta}+2\tilde{q}_{4}s_{2}\right\}$
$=\max\left\{\frac{2\tilde{q}_{2}\tau}{\rho\gamma|\tilde{\lambda}_{max}|},\frac{2(\tilde{q}_{3}+2(48\tau^{2}\gamma^{2}L^{2}\tilde{c}_{0}+\tilde{c}_{3})s_{2})}{\tau\rho\gamma|\tilde{\lambda}_{max}|}\right\}.$
Letting $\gamma\leq\min\{\bar{\gamma}_{1},\bar{\gamma}_{14}\}$, then
$\tilde{\delta}=\delta+\frac{\tilde{q}_{1}}{1-\delta}+2\tilde{q}_{4}(s_{0}+s_{1})<1.$ (65)
@@
$\left(1-\frac{\tilde{R}}{1-\tilde{\delta}}\frac{8}{\gamma\tau}\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)\right)\sum_{k=0}^{K-1}\mathcal{D}_{k}$
$\leq\frac{8\tilde{F}\left(\bar{x}_{0}\right)}{\gamma\tau}+\frac{8\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)}{\gamma\tau(1-\tilde{\delta})}\|\widehat{\mathbf{d}}_{0}\|^{2}.$
Similarly to the previous analysis, when $\gamma\leq\frac{\gamma_{14}}{2}=\bar{\gamma}_{15}$, $1-\tilde{\delta}\geq\frac{|\tilde{\lambda}_{\max}|\rho\tau\gamma}{4}.$ Let $\gamma\leq\min\{\bar{\gamma}_{16},\bar{\gamma}_{17}\}$, then
$\frac{\tilde{R}}{1-\tilde{\delta}}\frac{8}{\gamma\tau}\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)\leq\frac{1}{2},$
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\frac{16\tilde{F}\left(\bar{x}_{0}\right)}{K\gamma\tau}$
$+\frac{16}{K\gamma\tau}\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)\frac{1}{1-\tilde{\delta}}\|\widehat{\mathbf{d}}_{0}\|^{2}.$
By collecting all step-size conditions, if the step-size $\gamma$ satisfies (51), (56), (57), (59), (61), (65) and (C-B), then the states $\{\mathbf{X}_{k}\}$ generated by LT-ADMM-VR converge to the stationary point, concluding the proof.

### v2

@@ 9.1 Key bounds
$\tilde{\beta}_{0}$ $\coloneqq\frac{72\beta\tau^{2}}{\lambda_{l}\rho}+144\tau^{3}\beta^{2},$
$\tilde{c}_{0}$ $\coloneqq\gamma^{2}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})6\tau L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}+6L^{2}\frac{\gamma^{4}}{\beta^{2}}2\tau L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2},$
$\tilde{c}_{1}$ $\coloneqq\gamma^{2}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})4\tau L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}+6L^{2}\frac{\gamma^{4}}{\beta^{2}}2\tau L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2},$
$\tilde{c}_{2}$ $\coloneqq 6L^{2}\frac{\gamma^{4}}{\beta^{2}}N\|\widehat{\mathbf{V}}^{-1}\|^{2},$
$\tilde{q}_{0}$ $\coloneqq\tilde{c}_{0}\tilde{\beta}_{0}+2(s_{0}+s_{1})(\tilde{c}_{1}+32\tau^{2}\gamma^{2}L^{2}\tilde{c}_{0}),$
$\tilde{q}_{1}$ $\coloneqq 6L^{2}\frac{\gamma^{4}}{\beta^{2}}N,$
$\tilde{q}_{2}$ $\coloneqq 16\tau^{3}\gamma^{2}N\tilde{c}_{0}+2s_{2}(\tilde{c}_{1}+32\tau^{2}\gamma^{2}L^{2}\tilde{c}_{0}).$
Proof 9.13.
When $\beta\tau\leq\frac{2}{\lambda_{u}\rho}$ and $\gamma<\min\{\bar{\gamma}_{1},\bar{\gamma}_{7}\}$, substituting (49) and (50) into (28) then yields
$\|{\mathbf{h}}_{k}\|^{2}\leq\gamma^{2}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})(6\tau L^{2}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+4\tau L^{2}\mathbb{E}[{t}_{k}])$
$+6L^{2}\frac{\gamma^{4}}{\beta^{2}}\bigg(4\tau L^{2}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+4\tau L^{2}\mathbb{E}[{t}_{k}]+N\|\sum_{t=0}^{\tau-1}\overline{\nabla\mathrm{F}}(\mathrm{\Phi}_{k}^{t})\|^{2}\bigg)$
and
$\|\widehat{\mathbf{h}}_{k}\|^{2}\leq\tilde{c}_{0}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\tilde{c}_{1}t_{k}+\tilde{c}_{2}\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}$
$\leq\tilde{q}_{0}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+\tilde{q}_{1}\mathbb{E}[\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}]+\tilde{q}_{2}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}],$
together with (17), it proves that (61) holds.
9.2 Theorem 2
Based on (30), substituting $y=\bar{x}_{k+1}$ and $z=\bar{x}_{k}$ into (38) and using (49), we get
@@ 9.2 Theorem 2
$\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\gamma\mathbb{E}[\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}\sum_{t}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\rangle]$
$+\frac{\gamma^{2}L}{2}\mathbb{E}[\|\frac{1}{N}\sum_{t}\sum_{i}g_{i}(\phi_{i,k}^{t})\|^{2}]$
$\leq\mathbb{E}[F(\bar{x}_{k})]-\gamma[\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}\sum_{t}\sum_{i}\nabla f_{i}(\phi_{i,k}^{t}))\rangle]$
$+\gamma^{2}\tau L\mathbb{E}[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]+2\gamma^{2}\tau L^{3}(\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\mathbb{E}[{t}_{k}]).$
Using now $2\langle a,b\rangle=\|a\|^{2}+\|b\|^{2}-\|a-b\|^{2}$, we have
$-\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}\sum_{t}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\rangle$
$\leq-\frac{\tau}{2}\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}-\frac{1}{2}\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}+\frac{L^{2}}{2N}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}.$
Now, combining the two equations above and using (16), yields
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\frac{\gamma\tau}{2}\mathbb{E}[\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}]$
$-\frac{\gamma}{2}(1-2\gamma\tau L)\sum_{t}\mathbb{E}[\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\frac{\gamma L^{2}}{2N}\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]+2\gamma^{2}\tau L^{3}(\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\mathbb{E}[{t}_{k}]).$
Using (50) and (53) we have
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\frac{\gamma\tau}{2}\mathbb{E}[\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}]$
$-\frac{\gamma}{2}(1-2\gamma\tau L)\mathbb{E}[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\tilde{q}_{3}\mathbb{E}[\|\nabla F(x)\|^{2}]+\tilde{q}_{4}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}],$
@@
$\tilde{q}_{3}$ $\coloneqq 16\tau^{3}\gamma^{2}N(\frac{\gamma L^{2}}{2N}+2\gamma^{2}\tau L^{3})$
$+2s_{2}(2\gamma^{2}\tau L^{3}+32\tau^{2}\gamma^{2}L^{2}(\frac{\gamma L^{2}}{2N}+2\gamma^{2}\tau L^{3})),$
$\tilde{q}_{4}$ $\coloneqq(\frac{\gamma L^{2}}{2N}+2\gamma^{2}\tau L^{3})\tilde{\beta}_{0}$
$+2(s_{0}+s_{1})(2\gamma^{2}\tau L^{3}+32\tau^{2}\gamma^{2}L^{2}(\frac{\gamma L^{2}}{2N}+2\gamma^{2}\tau L^{3})).$
Letting $\gamma\leq\min\{1,\bar{\gamma}_{11},\bar{\gamma}_{12}\}$, then
$\tilde{q}_{3}\leq\frac{3\gamma\tau}{8},\quad 2\gamma\tau L\leq\frac{3}{4},$ (62)
@@
$-\frac{\gamma}{8}\mathbb{E}[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]+\tilde{q}_{4}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}].$
@@
$\mathcal{D}_{k}$ $\leq\frac{8}{\gamma\tau}(\mathbb{E}[\tilde{F}\left(\bar{x}_{k}\right)]-\mathbb{E}[\tilde{F}\left(\bar{x}_{k+1}\right)])+\frac{8}{\gamma\tau}\tilde{q}_{4}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}],$ (63)
where $\mathcal{D}_{k}$ is defined in (7), and $\tilde{F}\left(\bar{x}_{k}\right)=F\left(\bar{x}_{k}\right)-F(x^{*})$. Summing (63) over $k=0,1,\ldots,K-1$ and using $\tilde{F}\left(\bar{x}_{k}\right)\leq 0$, it holds that
$\sum_{r=0}^{K-1}\mathcal{D}_{k}\leq\frac{8\tilde{F}\left(\bar{x}_{0}\right)}{\gamma\tau}+\frac{8\tilde{q}_{4}}{\gamma\tau}\sum_{k=0}^{K-1}\|\mathbb{E}[\hat{\mathbf{d}}_{k}\|^{2}].$ (64)
According to (61), we derive that $\forall k\geq 0$,
$\mathbb{E}[\|\widehat{\mathbf{d}}_{k+1}\|^{2}]\leq(\delta+\frac{\tilde{q}_{0}}{1-\delta})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$ (65)
$+\frac{\tilde{q}_{1}}{1-\delta}\mathbb{E}[\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}]+\frac{\tilde{q}_{2}}{1-\delta}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]$
@@
$\tilde{R}=\max\left\{\frac{\tilde{q}_{1}\tau^{2}}{1-\delta},\frac{\tilde{q}_{2}}{1-\delta}\right\}.$
Letting $\gamma\leq\min\{\bar{\gamma}_{1},\bar{\gamma}_{13}\}$ and $\frac{1}{\tau\lambda_{u}\rho}\leq\beta<\frac{2}{\tau\lambda_{u}\rho}$, then
$\tilde{\delta}=\delta+\frac{\tilde{q}_{0}}{1-\delta}\leq 1-\frac{\lambda_{l}}{4\lambda_{u}}.$ (66)
@@
$\sum_{r=0}^{K-1}\mathcal{D}_{k}\leq\frac{8\tilde{F}\left(\bar{x}_{0}\right)}{\gamma\tau}+\frac{8\tilde{q}_{4}}{\gamma\tau}\sum_{k=0}^{K-1}\|\mathbb{E}[\hat{\mathbf{d}}_{k}\|^{2}].$ (68)
$(1-\frac{\tilde{R}}{1-\tilde{\delta}}\frac{8\tilde{q}_{4}}{\gamma\tau})\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\frac{8\tilde{F}\left(\bar{x}_{0}\right)}{\gamma\tau}+\frac{8\tilde{q}_{4}}{\gamma\tau(1-\tilde{\delta})}\|\widehat{\mathbf{d}}_{0}\|^{2}.$
Since $1-\tilde{\delta}\geq\frac{\lambda_{l}}{4\lambda_{u}},$ let $\gamma\leq\min\{\bar{\gamma}_{14},\bar{\gamma}_{15}\}$, then
$\frac{\tilde{R}}{1-\tilde{\delta}}\frac{8\tilde{q}_{4}}{\gamma\tau}\leq\frac{1}{2},$ (69)
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\frac{16\tilde{F}\left(\bar{x}_{0}\right)}{K\gamma\tau}+\frac{16\tilde{q}_{4}}{K\gamma\tau(1-\tilde{\delta})}\|\widehat{\mathbf{d}}_{0}\|^{2}.$ (70)
By collecting all step-size conditions, if the step-size $\gamma$ satisfies $\gamma\leq\min\bar{\gamma}_{i=1,7,8,\ldots,15}$, then the states $\{\mathbf{X}_{k}\}$ generated by LT-ADMM-VR converge to the stationary point, concluding the proof.